OpenStudy (anonymous):

Given that f(x) = x2 − 3x + 3 and g(x) = the quantity of x minus one, over four, solve for f(g(x)) when x = 5 How would I solve this problem?

zepdrix (zepdrix):

\[\Large \color{royalblue}{\text{Welcome To OpenStudy! :)}}\]

zepdrix (zepdrix):

\[\Large f(x)=x^2-3x+3 \qquad\qquad g(x)=\frac{x-1}{4}\] \[\Large g(\color{royalblue}{5})=\frac{\color{royalblue}{5}-1}{4} \qquad\to\qquad \color{#CC0033}{g(5)=1}\] \[\Large f\left(\color{#CC0033}{g(5)}\right) = f(\color{#CC0033}{1})\]

zepdrix (zepdrix):

\[\Large f(\color{#F35633}{x})=\color{#F35633}{x}^2-3\color{#F35633}{x}+3\]Understand how to find f(1) from here?\[\Large f(\color{#F35633}{1})=?\]

OpenStudy (anonymous):

I don't really follow.

zepdrix (zepdrix):

Hmm function notation can be a little tricky. Maybe there is another way to explain it. hmmmm

OpenStudy (anonymous):

I just really need an answer to make me understand. I've been relearning the function thing since last week and I'm having I problem learning it online.